OpenStudy (anonymous):

Solve the system of equations by either the addition method or the substitution method. Show your work. Check your solution. y = –3x + 11 –6x + 6y = –6

OpenStudy (texaschic101):

Well, you already know what y equals...it equals -3x + 11. Now sub -3x + 11 in for y in the second equation... -6x + 6y = -6 -6x + 6(-3x + 11) = - 6 Distribute the 6 through the parenthesis. Then combine like terms. Next, get x on one side and everything else on the other side. Then solve for x. Can you take it from here ?